### v1

@@ Abstract
Recently, large language and vision models have shown strong performance, but due to high pre-training and fine-tuning costs, research has shifted towards faster training via dataset pruning. Previous methods used sample loss as an evaluation criterion, aiming to select the most "difficult" samples for training. However, when the pruning rate increases, the number of times each sample is trained becomes more evenly distributed, which causes many critical or general samples to not be effectively fitted. We refer to this as Low-Frequency Learning (LFL). In other words, LFL prevents the model from remembering most samples. In our work, we decompose the scoring function of LFL, provide a theoretical explanation for the inefficiency of LFL, and propose adding a memory term to the scoring function to enhance the model’s memory capability, along with an approximation of this memory term. Similarly, we explore memory in Self-Supervised Learning (SSL), marking the first discussion on SSL memory. Using contrastive learning, we derive the memory term both theoretically and experimentally. Finally, we propose Enhance Memory Pruning (EMP), which addresses the issue of insufficient memory under high pruning rates by enhancing the model’s memory of data, thereby improving its performance. We evaluated the performance of EMP in tasks such as image classification, natural language understanding, and model pre-training. [^1] Our code was shown in https://github.com/xiaojinying/EMP The results show that EMP can improve model performance under extreme pruning rates. For example, in the CIFAR100-ResNet50 pre-training task, with 70% pruning, EMP outperforms current methods by 2.2%.
@@
[…]
{highlights}
Firstly, through theoretical analysis, we investigated the inefficiency of previous data pruning methods and their causes, which are referred to here as LFL (Low-Frequency Learning). In this work, we decomposed the scoring function to isolate a memory term that is negatively correlated with the scoring function. By analyzing the dynamic relationship between LFL and this memory term, we theoretically concluded that LFL leads to insufficient model memory of samples. In contrast, repeatedly learning from typical or general samples is our main research objective.
We have studied model memory and advocate for the addition of a memory term to the scoring function to enhance the model’s memory retention of samples. In this work, we specifically discuss two scenarios: Supervised Learning (SL) and Self-Supervised Learning (SSL). To our knowledge, this is the first time memory in SSL has been discussed. In SL, we use the mutual information term derived from the scoring function as the memory term, which quantifies the distribution relationship between model weights and samples. Since the mutual information term involves complex distributions of data and parameters, it is not easy to compute, and we have approximated it. In SSL, we focus on Contrastive Learning (CL), where we transfer the ideas from previous memory research in SL to CL. As the model’s memory of samples is specifically reflected in the hidden layers of the basic encoder, we establish a memory term through these hidden layers.
Based on the analysis above, we propose the Enhance Memory Pruning (EMP) method, which is a dynamic data pruning approach. We have evaluated the performance of EMP on image classification tasks, natural language understanding tasks, and model pre-training tasks. Experimental validation has shown that EMP leads the state-of-the-art methods in most cases and achieves commendable performance under high pruning rates. For instance, in image classification, with CIFAR100-ResNet18, pruning 70% of the dataset still leads other methods by at least 2.1%.
@@ 1 INTRODUCTION
Recently, deep learning technology has achieved significant breakthroughs in fields such as visual recognition (Shamshad et al., 2023; Radford et al., 2021; Chen et al., 2023) and natural language processing (Achiam et al., 2023; Devlin, 2018; Touvron et al., 2023). Although these models exhibit exceptional performance, they often require training and fine-tuning on large datasets, especially in the pre-training of large language models (LLMs). This process not only demands substantial computational resources but also consumes a considerable amount of time. Therefore, reducing the burden of training models on large-scale datasets has become increasingly important for promoting the application of deep learning technology across a broader range of fields.
Data pruning accelerates model pre-training by retaining a core subset of typical or general samples, thereby reducing the number of training iterations. Recently, dynamic pruning methods that adjust the retained dataset in real-time have become popular (Qin et al., ; Raju et al., 2021; Xiao et al., 2024). These methods dynamically score data at each checkpoint, allowing pruned data to be retrained. Since these methods do not require additional training of a proxy network, they have started to gain popularity. Some of these methods use sample loss to construct (Qin et al., ; Raju et al., 2021), selecting the most "difficult" samples to train at each checkpoint. Raju et al. (2021) indicates that such methods can acquire more information and knowledge. However, we found that these methods are only effective at specific pruning rates. In other words, when the pruning rate is high, these methods exhibit poor performance. Moreover, on datasets that are difficult to fit, such as ImageNet-1K, the performance of low-frequency learning is comparable to, or even lower than, dynamic random pruning methods (see Table 2).
Naturally, we explored the reasons behind this. It is worth noting that these methods retain samples with the highest loss at each checkpoint. However, during the training process, due to the model’s strong fitting ability (Zhang et al., 2021), the loss of these samples will eventually decrease (Amari, 1993), and these samples are often pruned in the next selection. This results in an even distribution of the number of times samples are selected, which we refer to as Low-Frequency Learning (LFL). As shown in Figure (1), we counted the number of times each sample was selected during the entire training process and found that in LFL, the selection frequency of samples is more concentrated. In other words, the difference in the number of times each sample is included in the training is small, and the number of times is generally low. This limits the number of times each sample is trained. Previous work has shown that repeated learning can enhance model memory (Wei et al., 2024); for instance, in language models, the model tends to capture familiar phrases and commonly accepted knowledge (Zhang et al., 2023).
@@ 2 RELATED WORK
Static Pruning. Static pruning aims to select a compact core subset before training. Toneva et al. analyzed the forgetting of samples during the training process and assessed their forgetfulness in training to eliminate unforgettable samples, Xia et al. (2022) selected a moderate core set based on the distance of data points to the center, Aljundi et al. (2019) utilized the maximum diversity of samples in the gradient space for selection, Yang et al. ingeniously used the influence function, constructing a core subset based on the impact of samples on the model’s generalization ability. Tan et al. (2024) designed a first-order gradient approximation to assess the impact of samples on the optimal empirical risk. Although these methods select an efficient core subset, the core subset, as a form of a posteriori knowledge, must be learned through the optimization of one or more proxy models to understand the characteristics and distribution of the data. For example, in Yang et al. , SENet and ResNet18 are used as proxy models to accelerate the training of ResNet50. Although these proxy models are smaller in scale, they require training on the entire dataset, which inevitably brings additional overhead. More importantly, these methods result in a core subset with poor generalizability. For instance, in Yang et al. , for different specifications of ResNet, the proxy model also requires additional selection and optimization.
Dynamic Pruning. To address the aforementioned additional overhead, Raju et al. (2021) argued that the optimal dynamic scoring is closely integrated with the training trajectory of the model. They categorized data into three types based on the number of times samples were selected and found that these samples are highly variable and can transition within the training dynamics. They first performed dataset pruning on the retained data’s loss at each checkpoint during training without a proxy network, and from their conclusions, dynamic pruning is always superior to static pruning, even random selection. In Qin et al. , soft pruning was proposed; they believed that for a dataset with N samples, hard pruning with an unchanged pruning rate requires a complexity of $\mathcal{O}(\log N)$ for sorting the scores, while soft pruning only requires $\mathcal{O}(1)$. Although the sorting cost is reduced, experimentally, soft pruning cannot determine the true pruning ratio, thereby explicitly increasing the training cost of the model. It is worth noting that before us, Qin et al. ; Raju et al. (2021) provided state-of-the-art performance in dynamic pruning.
@@ 3.1 Problem Description
Given a large-scale dataset $D=\{X,Y\}$ containing $n$ training samples, where the input $X=\{x^{(1)},\ldots,x^{(n)}\}$, labels $Y=\{y^{(1)},\ldots,y^{(n)}\}$, and $f(Y|X,\theta)$ represents the network output parameterized by $\theta$. The goal of the data is to identify a subset $\hat{D}=\{\hat{X},\hat{Y}\}$, where $\hat{D}\subseteq D$, thereby accelerating the model training process. It is important to note that the pruning rate $s$ is expressed as $s=\frac{\|D-D_{0}\|_{0}}{\|D_{0}\|_{0}}$. In this paper, the data is generated by the distribution $p(x,y)$, and we need to use several information quantities such as entropy: $H(X)=-\mathbb{E}[\log p(x)]$, mutual information: $I(X;Y)=H(X)+H(Y)-H(X,Y)$, and Kullback-Leibler divergence: $KL(p(x)||q(x))=\mathbb{E}_{x\sim p(x)}[\log(p(x)/q(x))]$.
@@ 3.2 LFL Leads to Poor Memory
Previous works (Qin et al., ; Raju et al., 2021) used loss values to score samples, and at each checkpoint, they selected the portion with the highest loss values according to the pruning rate. We believe this is a form of low frequency learning. Specifically, at the first checkpoint, they selected $\hat{D}_{1}$, and due to the model’s gradient descent on $L(\hat{D}_{1},\theta)$ during the training process, these data are very likely not to be retained at the next checkpoint. On the contrary, the algorithm tends to retain the data in $D_{-\hat{D}_{1}}$, at which point the chance of data being retained is averaged, as shown in Figure (1). However, repeated learning can strengthen the model’s memory (Wei et al., 2024), for example, in language models, the model is more inclined to capture familiar phrases, recognized knowledge (Zhang et al., 2023). Intuitively, LFL does not provide the model with the opportunity for repeated learning, leading to insufficient memory of the data by the model, especially under high pruning rates.
In our experiments, we tested the aforementioned viewpoint, as shown in Figure (3), where we established LFL under an extreme condition where each sample is trained the same number of times, which we refer to as Extreme Low-Frequency Learning (ELFL). It is not difficult to see from the graph that under high pruning rates, the model’s training accuracy is not high. This situation is not unexpected, as it can be observed that in the early stages of training, due to the inconsistent gradient information from atypical and noisy examples, they may cancel each other out (Stephenson et al., ; Gu and Tresp, 2019), at which point the model tends to learn general patterns (Wei et al., 2024), and the rise in training accuracy is not significant. In the later stages of training, the InfoBatch method, which uses loss as a criterion, and ELFL do not show a significant increase in training accuracy, which is particularly prominent in the challenging CIFAR100 dataset. Combining this with the long-tail theory (Feldman, 2020; Feldman and Zhang, 2020), we believe that in the later stages, important or typical samples appear infrequently, and these samples require more repeated memorization, while the model only remembers those long-tail samples that are easier to remember. Therefore, based on the experiments, pruning methods that use loss as a criterion struggle to memorize the data.
@@
In previous work (Harutyunyan et al., 2020; Achille and Soatto, 2018), when the data is generated consistently, the training loss (scoring function) in the form of cross-entropy can be decomposed as:
@@
However, it can be noted that, in order to obtain more information (Toneva et al., ), previous data pruning methods (Toneva et al., ; Qin et al., ; Raju et al., 2021) select the samples with the highest training loss at each checkpoint. In fact, such algorithms can achieve this by minimizing the second term in Equation (1).
[…]
@@
For a set of $m$ samples, an independently and identically distributed subset of retained data $(\hat{X},\hat{Y})=\{(x^{(1)},y^{(1)}),\ldots,(x^{(m)},y^{(m)})\}$, let $\hat{y}^{(i)}$ represent the model’s prediction for the $i$-th sample, and let $c^{(i)}=\mathbb{1}\{\hat{y}^{(i)}=y^{(i)}\}$ be the correctness variable corresponding to the prediction $\hat{y}^{(i)}$. Thus, the following inequality holds:
@@
Under this Markov chain, Fano’s inequality (Fano, 2008) provides a lower bound on the training error:
@@
For the retained dataset $(\hat{X},\hat{Y})$ with a sample size of $m$, the sum can be obtained as follows:
@@
Theorem 1 establishes an expected upper bound on the training accuracy on the retained dataset, and it can be observed that this upper bound decreases as $I(\theta;Y|X)$ decreases. As mentioned earlier, by selecting for learning based on the largest losses, such algorithms can achieve this by minimizing the second term in Equation (1). In this case, the $I(\theta;\hat{Y}|\hat{X})$ of the retained data tends to decrease, and according to Equation (2), we can intuitively consider that the training accuracy will decrease at this time.
@@
$\forall z_{i}\in D,\forall z_{j}\in A\left(D\right)_{k},\mathcal{L}\left(z_{i},\theta_{k}\right)\leq\mathcal{L}\left(z_{j},\theta_{k}\right)$ Then the resulting model has a weak memory of the data, in other words, the model fits the training data poorly.
@@
According to Equation (1), when fitting the model with the subset of data that has the highest training loss, algorithms with limited memory capacity for labels will cause $I(\theta;\hat{Y}|\hat{X})$ to tend towards 0. At this point, according to Theorem 1, this will reduce the upper bound of training accuracy, thereby affecting the model’s fit.
@@ 3.3 Enhancing Memory in SL
In Section 3.2, the inability of LFL to enable model memorization of data was analyzed. In the context of supervised learning, looking at Equation (1), we need to increase $I(\theta;Y|X)$ to enable the model to remember the data. To achieve this goal, it is first necessary to estimate $I(\theta;Y|X)$, which involves the independence of $\theta$ and $Y$. Previous work has used gradients (Harutyunyan et al., 2020) and differential methods (Sordoni et al., 2021) to estimate mutual information. In our work, we decompose $I(\theta;Y|X)$ as follows:
@@
In which, $\theta_{k}$ represents the current model parameters, $\mathcal{L}(\bullet)$ generally refers to the cross-entropy function, and $\beta$ is an adjustable hyperparameter that balances the model’s learning of general patterns and memorization of data. The core idea of this scoring function is to use the sample loss to enable the model to obtain as much information as possible, allowing the model to learn not only general patterns but also, through the second term, to prefer samples that are easy to remember, achieving higher model performance.
@@
For a subset $\hat{D}\subseteq D$, if $D_{k}$ satisfies:
@@
We assume that the subset generated by LFL is $\hat{D}_{\text{LFL}}$. Previous work (Koltchinskii and Panchenko, 2002) has expressed the upper bound on generalization error, which in our context can be represented as:
@@
Where $\mathcal{R}(\hat{\theta}_{\hat{D}})$ is the expected loss, ^$\hat{\mathcal{R}}(\hat{D},\hat{\theta}_{\hat{D}})$ is the empirical risk on $\hat{D}$, which is the fitting loss, $\varepsilon$ is a coefficient related to the model size and the size of the retained dataset, which remains fixed after the pruning rate and model architecture are set, and $\hat{\theta}_{\hat{D}}$ represents the optimal parameters obtained after optimization with $\hat{D}$. Our method enhances the model’s memory, and likewise, improves the model’s fitting capability. Therefore, based on the above experimental and theoretical analysis, we can conclude that:
^$\hat{\mathcal{R}}\left(\hat{D},{\hat{\theta}}_{\hat{D}}\right)<\hat{\mathcal{R}}\left({\hat{D}}_{\text{LFL}},{\hat{\theta}}_{\hat{D}_{\text{LFL}}}\right)$ (10)
@@
0:  Dataset $D$, Epoch $T$, Pruning Ratio $s$, Initial Model $\theta$, Retained Dataset $\hat{D}_{0}=D$.
1:  for $t$ from 0 to $T-1$ do
2:     if $t!=0$ then
3:        Select the retained dataset $\hat{D}_{t}$ based on the pruning rate $s$ and sample scores.
4:     else
5:        $\hat{D}_{t}=D$
6:     end if
7:     Get train sequence $\hat{D}_{t}=\{B_{0},B_{1},,,B_{b_{t}-1}\}$
8:     for $j$ from 0 to $b_{t}-1$ do
9:        Get loss $L(B_{j},\theta_{t}^{j})$
10:        Update score with Equation (8)
11:        Update model with optimizer
12:     end for

### v2

@@ Abstract
Large language and vision models have demonstrated remarkable performance, but their high pre-training and fine-tuning costs have led to growing interest in accelerating training through dataset pruning. Traditional pruning approaches typically rely on sample loss to identify and retain the most "difficult" samples. However, as the pruning rate increases, the training frequency of each sample becomes more uniform, which can result in the underfitting of critical or general samples. We identify this phenomenon as Low-Frequency Learning (LFL), where the model fails to retain knowledge of the majority of samples. In this work, we decompose the scoring function of LFL, providing a theoretical analysis of its inefficiencies. To counteract this issue, we propose a novel enhancement to the scoring function by introducing a memory term designed to strengthen the model’s ability to memorize essential data. We also offer an approximation for this memory term. Additionally, we extend this concept to Self-Supervised Learning (SSL), marking the first investigation into the role of memory in SSL. By leveraging contrastive learning, we derive the memory term both theoretically and experimentally. Based on these insights, we introduce Enhance Memory Pruning (EMP), a technique that mitigates memory loss under high pruning rates by improving the model’s data retention. EMP has been evaluated across various tasks, including image classification, natural language understanding, and model pre-training. Our experiments demonstrate that EMP significantly enhances model performance, particularly under extreme pruning conditions. For instance, in the CIFAR100-ResNet50 pre-training task with a 70% pruning rate, EMP outperforms current state-of-the-art methods by 2.2%. [^1] The code was shown in https://github.com/xiaojinying/EMP
@@
Data Pruning, Model Memory, Mutual Information, Self-Supervised Learning, Supervised Learning
@@ 1 Introduction
Recently, deep learning technology has achieved significant breakthroughs in fields such as visual recognition [ref20, ref21, ref22] and natural language processing [ref23, ref24, ref25]. Although these models exhibit exceptional performance, they often require training and fine-tuning on large datasets, especially in the pre-training of large language models (LLMs). This process not only demands substantial computational resources but also consumes a considerable amount of time. Therefore, reducing the burden of training models on large-scale datasets has become increasingly important for promoting the application of deep learning technology across a broader range of fields.
Data pruning accelerates model pre-training by retaining a core subset of typical or general samples, thereby reducing the number of training iterations. Recently, dynamic pruning methods that adjust the retained dataset in real-time have become popular [ref4, ref5, ref27]. These methods dynamically score data at each checkpoint, allowing pruned data to be retrained. Since these methods do not require additional training of a proxy network, they have started to gain popularity. Some of these methods use sample loss to construct [ref4, ref5], selecting the most "difficult" samples to train at each checkpoint. [ref5] indicates that such methods can acquire more information and knowledge. However, we found that these methods are only effective at specific pruning rates. In other words, when the pruning rate is high, these methods exhibit poor performance. Moreover, on datasets that are difficult to fit, such as ImageNet-1K, the performance of low-frequency learning is comparable to, or even lower than, dynamic random pruning methods (see Table 2).
Naturally, we explored the reasons behind this. It is worth noting that these methods retain samples with the highest loss at each checkpoint. However, during the training process, due to the model’s strong fitting ability [ref28], the loss of these samples will eventually decrease [ref26], and these samples are often pruned in the next selection. This results in an even distribution of the number of times samples are selected, which we refer to as Low-Frequency Learning (LFL). As shown in Figure (1), we counted the number of times each sample was selected during the entire training process and found that in LFL, the selection frequency of samples is more concentrated. In other words, the difference in the number of times each sample is included in the training is small, and the number of times is generally low. This limits the number of times each sample is trained. Previous work has shown that repeated learning can enhance model memory [ref18]; for instance, in language models, the model tends to capture familiar phrases and commonly accepted knowledge [ref8].
@@ 2 Related Work
Static Pruning. Static pruning aims to select a compact core subset before training. [ref3] analyzed the forgetting of samples during the training process and assessed their forgetfulness in training to eliminate unforgettable samples, [ref44] selected a moderate core set based on the distance of data points to the center, [ref43] utilized the maximum diversity of samples in the gradient space for selection, [ref33] ingeniously used the influence function, constructing a core subset based on the impact of samples on the model’s generalization ability. [ref17] designed a first-order gradient approximation to assess the impact of samples on the optimal empirical risk. Although these methods select an efficient core subset, the core subset, as a form of a posteriori knowledge, must be learned through the optimization of one or more proxy models to understand the characteristics and distribution of the data. For example, in [ref33], SENet and ResNet18 are used as proxy models to accelerate the training of ResNet50. Although these proxy models are smaller in scale, they require training on the entire dataset, which inevitably brings additional overhead. More importantly, these methods result in a core subset with poor generalizability. For instance, in [ref33], for different specifications of ResNet, the proxy model also requires additional selection and optimization.
Dynamic Pruning. To address the aforementioned additional overhead, [ref5] argued that the optimal dynamic scoring is closely integrated with the training trajectory of the model. They categorized data into three types based on the number of times samples were selected and found that these samples are highly variable and can transition within the training dynamics. They first performed dataset pruning on the retained data’s loss at each checkpoint during training without a proxy network, and from their conclusions, dynamic pruning is always superior to static pruning, even random selection. In [ref4], soft pruning was proposed; they believed that for a dataset with N samples, hard pruning with an unchanged pruning rate requires a complexity of $\mathcal{O}(\log N)$ for sorting the scores, while soft pruning only requires $\mathcal{O}(1)$. Although the sorting cost is reduced, experimentally, soft pruning cannot determine the true pruning ratio, thereby explicitly increasing the training cost of the model. It is worth noting that before us, [ref4, ref5] provided state-of-the-art performance in dynamic pruning.
@@ 3.1 Problem Description
Given a large-scale dataset $D=\{X,Y\}$ containing $n$ training samples, where the input $X=\{x^{(1)},\ldots,x^{(n)}\}$, labels $Y=\{y^{(1)},\ldots,y^{(n)}\}$, and $f(Y|X,\theta)$ represents the network output parameterized by $\theta$. The goal of the data is to identify a subset $\hat{D}=\{\hat{X},\hat{Y}\}$, where $\hat{D}\subseteq D$, thereby accelerating the model training process. It is important to note that the pruning rate $s$ is expressed as $s=\frac{\|D-D_{0}\|_{0}}{\|D_{0}\|_{0}}$. In this paper, the data is generated by the distribution $p(x,y)$, and we need to use several information quantities such as entropy: $H(X)=-\mathbb{E}[\log p(x)]$, mutual information: I(X;Y)=H(X)+H(Y)H(X,Y)I(X;Y)=H(X)+H(Y)-H(X,Y), and Kullback-Leibler divergence: $KL(p(x)||q(x))=\mathbb{E}_{x\sim p(x)}[\log(p(x)/q(x))]$.
@@ 3.2 LFL Leads to Poor Memory
Previous works [ref4, ref5] used loss values to score samples, and at each checkpoint, they selected the portion with the highest loss values according to the pruning rate. We believe this is a form of low frequency learning. Specifically, at the first checkpoint, they selected $\hat{D}_{1}$, and due to the model’s gradient descent on $L(\hat{D}_{1},\theta)$ during the training process, these data are very likely not to be retained at the next checkpoint. On the contrary, the algorithm tends to retain the data in $D_{-\hat{D}_{1}}$, at which point the chance of data being retained is averaged, as shown in Figure (1). However, repeated learning can strengthen the model’s memory [ref18], for example, in language models, the model is more inclined to capture familiar phrases, recognized knowledge [ref8]. Intuitively, LFL does not provide the model with the opportunity for repeated learning, leading to insufficient memory of the data by the model, especially under high pruning rates.
In our experiments, we tested the aforementioned viewpoint, as shown in Figure (3), where we established LFL under an extreme condition where each sample is trained the same number of times, which we refer to as Extreme Low-Frequency Learning (ELFL). It is not difficult to see from the graph that under high pruning rates, the model’s training accuracy is not high. This situation is not unexpected, as it can be observed that in the early stages of training, due to the inconsistent gradient information from atypical and noisy examples, they may cancel each other out [ref9, ref10], at which point the model tends to learn general patterns [ref18], and the rise in training accuracy is not significant. In the later stages of training, the InfoBatch method, which uses loss as a criterion, and ELFL do not show a significant increase in training accuracy, which is particularly prominent in the challenging CIFAR100 dataset. Combining this with the long-tail theory [ref11, ref19], we believe that in the later stages, important or typical samples appear infrequently, and these samples require more repeated memorization, while the model only remembers those long-tail samples that are easier to remember. Therefore, based on the experiments, pruning methods that use loss as a criterion struggle to memorize the data.
@@
In previous work [ref1, ref2], when the data is generated consistently, the training loss (scoring function) in the form of cross-entropy can be decomposed as:
@@
However, it can be noted that, in order to obtain more information [ref3], previous data pruning methods [ref3, ref4, ref5] select the samples with the highest training loss at each checkpoint. In fact, such algorithms can achieve this by minimizing the second term in Equation (1).
@@
For a set of $m$ samples, an independently and identically distributed subset of retained data $(\hat{X},\hat{Y})=\{(x^{(1)},y^{(1)}),\ldots,(x^{(m)},y^{(m)})\}$, let $\hat{y}^{(i)}$ represent the model’s prediction for the $i$-th sample, and let $c^{(i)}=\mathbf{1}\{\hat{y}^{(i)}=y^{(i)}\}$ be the correctness variable corresponding to the prediction $\hat{y}^{(i)}$. Thus, the following inequality holds:
@@
Under this Markov chain, Fano’s inequality [ref7] provides a lower bound on the training error:
@@
[…]
@@
[…]
@@
$\forall z_{i}\in D,\forall z_{j}\in A\left(D\right)_{k},\mathcal{L}\left(z_{i},\theta_{k}\right)\leq\mathcal{L}\left(z_{j},\theta_{k}\right)$
Then the resulting model has a weak memory of the data, in other words, the model fits the training data poorly.
@@
[…]
@@ 3.3 Enhancing Memory in SL
In Section 3.2, the inability of LFL to enable model memorization of data was analyzed. In the context of supervised learning, looking at Equation (1), we need to increase $I(\theta;Y|X)$ to enable the model to remember the data. To achieve this goal, it is first necessary to estimate $I(\theta;Y|X)$, which involves the independence of $\theta$ and $Y$. Previous work has used gradients [ref1] and differential methods [ref12] to estimate mutual information. In our work, we decompose $I(\theta;Y|X)$ as follows:
@@
In which, $\theta_{k}$ represents the current model parameters, $\mathcal{L}(\cdot)$ generally refers to the cross-entropy function, and $\beta$ is an adjustable hyperparameter that balances the model’s learning of general patterns and memorization of data. The core idea of this scoring function is to use the sample loss to enable the model to obtain as much information as possible, allowing the model to learn not only general patterns but also, through the second term, to prefer samples that are easy to remember, achieving higher model performance.
@@
For a subset $\hat{D}\subseteq D$, if $\hat{D}$ satisfies:
@@
We assume that the subset generated by LFL is $\hat{D}_{\text{LFL}}$. Previous work [ref29] has expressed the upper bound on generalization error, which in our context can be represented as:
@@
[…]
$\hat{\mathcal{R}}\left(\hat{D},{\hat{\theta}}_{\hat{D}}\right)<\hat{\mathcal{R}}\left({\hat{D}}_{\text{LFL}},{\hat{\theta}}_{\hat{D}_{\text{LFL}}}\right)$ (10)
@@
1:Dataset $D$, Epoch $T$, Pruning Ratio $s$, Initial Model $\theta$, Retained Dataset $\hat{D}_{0}=D$.
2:for $t$ from 0 to T1T-1 do
3:  if $t!=0$ then
4:   Select the retained dataset $\hat{D}_{t}$ based on the pruning rate $s$ and sample scores.
5:  else
6:   $\hat{D}_{t}=D$
7:  end if
8:  Get train sequence $\hat{D}_{t}=\{B_{0},B_{1},,,B_{b_{t}-1}\}$
9:  for $j$ from 0 to $b_{t}-1$ do
10:   Get loss $L(B_{j},\theta_{t}^{j})$
11:   Update score with Equation (8)
12:   Update model with optimizer